OpenStudy (anonymous):

Which expression is equivalent to 33 • 36 • 24 ? A 26 • 35 B 28 • 36 C 26 • 36 D 28 • 35

OpenStudy (anonymous):

The prime factorization of \(33\times36\times24\) is \((3\times11)\times(2^2\times3^2)\times(2^3\times3)=2^5\times3^4\times11\). Get the prime factorization of each choice, and see whether it matches with the above.